Grade 12th passMechanics

A ball is thrown vertically upwards with certain velocity. While moving upwards it reaches a point A at t1 second and then reaches the ground at t1+ t2
second. Find the height of point A above point of projection?

farheenkhan
9 Years agoGrade 12th pass

1 Answer

Pushkar Prateek
9 Years ago
Let initial projected speed pf ball be ‘u’ vertically up.
So in time t1
h = ut1 – ½ gt2    ( we have to find h )......................................................................(1)
 
We ball returns to the ground the displacement is ZERO.
S = ut +1/2 gt2
0 = u(t​+ t2) – ½ g (t​+ t2)2
Solvinng quadratic,
t1 + t2  = 0  or 2u/g  (neglectinng zero as it is the initial time)
t​+ t2 = 2u/g...............................................(2)
 
 
Substitute ‘u’  from 2 in 1
h =1/2 gt1t2